Prove A=B If cosA=cosB .Can we put both the angles in the same right triangle?

Grade:10

1 Answers

Umakant biswal
7 years ago
@ juhi 
yes , u can take that . 
kindly follow the below method for solving such question 
TAKE A RIGHT TRIANGLE ABC , RIGHT ANGLED AT C 
HERE 
COS A = AC/AB 
COS B = BC / AB 
GIVEN COS A = COS B 
AC/AB = BC / AB 
SO, A=B 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..